The difference between the length and width of a rectangle is 8 units. The peremeter is 84 units. Find the length and width.
Answer by ikleyn(52752)

.
One solution

L + W = 42   (half of the perimeter, or 84/2)
L - W =  8
----------------------Add the two equations

2L = 42 + 8 

2L = 50  ====>  L = 50/2 = 25.

Then W = 42 - 25 = 17.


Answer.  The length is 25 units, the width is 17 units.


Another solution.

L + W = 42


L + (L-8) = 42

2L - 8 = 42

2L = 42 + 8 = 50


L = 50/2 = 25,  and you get the same answer.